OpenStudy (hannahduhe):

find the length of the missing side. leave your answer in simplest radical form. LOOK IN COMMENTS FOR DRAWING.

OpenStudy (hannahduhe):



OpenStudy (gniiij):

hi again

OpenStudy (gniiij):

I can help

OpenStudy (gniiij):

so you do 6^2+x^2=15^2

OpenStudy (gniiij):

it is \[\pm3\sqrt{21}\]

OpenStudy (hannahduhe):

thanks soo much @GNiiij